# Chinese Spelling Correction as Rephrasing Language Model

### 18 Dec 2023 v2

@@ Abstract
This paper studies Chinese Spelling Correction (CSC), which aims to detect and correct the potential spelling errors in a given sentence. Current state-of-the-art methods regard CSC as a sequence tagging task and fine-tune BERT-based models on sentence pairs. However, we note a critical flaw in the process of tagging one character to another, that the correction is excessively conditioned on the error. This is opposite from human mindset, where individuals rephrase the complete sentence based on its semantics, rather than solely on the error patterns memorized before. Such a counter-intuitive learning process results in the bottleneck of generalizability and transferability of machine spelling correction. To address this, we propose Rephrasing Language Model (ReLM), where the model is trained to rephrase the entire sentence by infilling additional slots, instead of character-to-character tagging. This novel training paradigm achieves the new state-of-the-art results across fine-tuned and zero-shot CSC benchmarks, outperforming previous counterparts by a large margin. Our method also learns transferable language representation when CSC is jointly trained with other tasks. [^1] https://github.com/Claude-Liu/ReLM [^2] https://github.com/gingasan/lemon
@@ Introduction
Our contributions are summarized below: (1) We propose ReLM to narrow the gap between machine spelling correction and human spelling correction. (2) ReLM significantly enhances the generalizability of CSC models and refreshes the new state-of-the-art results across fine-tuned and zero-shot CSC benchmarks. (3) We probe into and enhance the transferability of CSC to other tasks. (4) Our analysis shows that ReLM effectively exploits and retains the pre-trained knowledge within PLMs, while tagging models do not.
@@
ECSpell  ECSpell (Lv et al. 2022) is a CSC benchmark with three domains, law (LAW, 1,960 training and 500 test samples), medical treatment (MED, 3,000 training and 500 test samples), and official document writing (ODW, 1,728 training and 500 test samples).
@@
SIGHAN  SIGHAN (Tseng et al. 2015) is a traditional CSC benchmark collected from the Chinese essays written by foreign speakers. Following Wu et al. (2023c), we evaluate the model on SIGHAN as zero-shot learning (1,100 test samples).
@@
We train the multi-task model on three distinct tasks, ECSpell for CSC, AFQMC for semantic similarity, and TNEWS for news classification. The three datasets are mixed together and we uniformly sample one batch from them during training. For the tagging models, we train three task-specific linear classifiers and one shared encoder. For ReLM, we share the entirety of model parameters for all three tasks. For ReLM with prompt, we train an additional prompt embeddings for each task.
@@
We investigate two mask strategies of the auxiliary MLM when training ReLM. The first is to uniformly mask any characters in the sentence and the second is to mask non-error characters only. From Table 6, we can see that both mask strategies are effective, while masking non-error characters works better. This is because masking error characters can reduce the amount of the real errors within the training data, which the model needs for learning correction.
@@
We also investigate the impact of the mask rate. From Table 7, we find that the performance on ECSpell keeps improving when the mask rate grows from 0% to 30%, and 30% is the best choice, while it drops when the mask rate is above 30%. While it still relies on specific data, it shows that the mask rate for ReLM is higher than that for MLM (Devlin et al. 2019). This is because ReLM is essentially a further refinement of PLMs, which in turn allows a higher mask rate.
@@
The third case is quite puzzling (Judicial power is not execution, but judgemental power.), especially the first error. The correct answer not only necessitates the semantics but also a legal principle that “judicial power is judgemental power”, which can only be attained through the pre-training process. We find that the tagging model does not possess such expertise and its answer is “law enforcement power is judgemental power”. It suggests that ReLM effectively inherits the knowledge of PLMs, while the tagging model results in the forgetfulness of what it has learned even enhanced with masked-fine-tuning.

### 28 Feb 2024 v3

@@ Abstract
This paper studies Chinese Spelling Correction (CSC), which aims to detect and correct the potential spelling errors in a given sentence. Current state-of-the-art methods regard CSC as a sequence tagging task and fine-tune BERT-based models on sentence pairs. However, we note a critical flaw in the process of tagging one character to another, that the correction is excessively conditioned on the error. This is opposite from human mindset, where individuals rephrase the complete sentence based on its semantics, rather than solely on the error patterns memorized before. Such a counter-intuitive learning process results in the bottleneck of generalizability and transferability of machine spelling correction. To address this, we propose Rephrasing Language Model (ReLM), where the model is trained to rephrase the entire sentence by infilling additional slots, instead of character-to-character tagging. This novel training paradigm achieves the new state-of-the-art results across fine-tuned and zero-shot CSC benchmarks, outperforming previous counterparts by a large margin. Our method also learns transferable language representation when CSC is jointly trained with other tasks.
@@ Introduction
Our contributions are summarized below: (1) We propose ReLM to narrow the gap between machine spelling correction and human spelling correction. (2) ReLM significantly enhances the generalizability of CSC models and refreshes the new state-of-the-art results across fine-tuned and zero-shot CSC benchmarks. (3) We probe into and enhance the transferability of CSC to other tasks. (4) Our analysis shows that ReLM effectively exploits and retains the pre-trained knowledge within PLMs, while tagging models do not.  [^1] https://github.com/Claude-Liu/ReLM [^2] https://github.com/gingasan/lemon
@@
ECSpell  ECSpell (Lv et al. 2022) is a CSC benchmark with three domains, LAW (1,960 training and 500 test samples), MED (3,000 training and 500 test samples), and ODW (1,728 training and 500 test samples).
@@
SIGHAN  SIGHAN (Tseng et al. 2015) is a CSC benchmark collected from the Chinese essays written by foreign speakers. Following Wu et al. (2023c), we evaluate the model on SIGHAN as zero-shot learning.
@@
We train the multi-task model on three distinct tasks, ECSpell for CSC, AFQMC for semantic similarity, and TNEWS for news classification. The three datasets are mixed together and we uniformly sample one batch from them during training. We fine-tune each model on all tasks for 15 epochs, with the batch size selected from {32, 128} and learning rate from {2e-5, 5e-5}. For the tagging models, we train three task-specific linear classifiers and one shared encoder. For ReLM, we share the entirety of model parameters for all three tasks. For ReLM with prompt embeddings, we train an additional prompt embeddings for each task. The MFT technique is exclusively applied on CSC.
@@
We investigate two mask strategies of the auxiliary MLM when training ReLM. The first is to uniformly mask any characters in the sentence and the second is to mask non-error characters only. From Table 6, we can see that both mask strategies are effective, while masking non-error characters works better. This is because masking error characters can reduce the amount of the real errors within the training data, which the model needs for learning correction. Additionally, we find that adding learnable prompts can further improve the performance of ReLM. The results in Table 6 and Table 7 are based on ReLM with learnable prompts.
@@
We also investigate the impact of the mask rate. From Table 7, we find that the performance on ECSpell keeps improving when the mask rate grows from 0% to 30%, and 30% is the best choice. While it still relies on specific data, it shows that the mask rate for ReLM is higher than that for MLM (Devlin et al. 2019). This is because ReLM is essentially a further refinement of PLMs, which allows a higher mask rate.
@@
The third case is quite puzzling (Judicial power is not execution, but judgemental power.), especially the first error. The correct answer not only necessitates the semantics but also a legal principle that “judicial power is judgemental power”, which can only be attained through the pre-training process. We find that the tagging model does not possess such expertise and its answer is “law enforcement power is judgemental power”. It suggests that ReLM effectively inherits the knowledge of PLMs, while the tagging model does not even enhanced with masked-fine-tuning.
@@ Conclusion
Acknowledgements
This paper was partially supported by Joint Research Project of Yangtze River Delta Science and Technology Innovation Community (No. 2022CSJGG1400).